# Data Augmentation for Time-Series Classification: a Comprehensive Survey

### v4

@@ Abstract
Data Augmentation (DA) has emerged as an indispensable strategy in time series classification (TSC), mainly due to its ability to amplify training samples, thus bolstering model robustness, diversifying datasets, and curtailing overfitting. However, the current landscape of DA in TSC is plagued with fragmented literature reviews, nebulous methodological taxonomies, inadequate evaluative measures, and a dearth of accessible and user-oriented tools. In light of these challenges, this study embarks on an exhaustive dissection of DA methodologies within the TSC realm. Our initial approach involved an extensive literature review that spanned a decade, revealing that contemporary surveys scarcely capture the breadth of advancements in DA for TSC, requiring us to meticulously analyze more than 100 scholarly articles to distill more than 60 unique DA techniques. This rigorous analysis precipitated the formulation of a novel taxonomy, designed for the intricacies of DA in TSC, categorizing techniques into five principal echelons: Transformation-Based, Pattern-Based, Generative, Decomposition-Based, and Automated Data Augmentation. Our taxonomy promises to serve as a robust navigational aid for scholars, offering clarity and direction in method selection. Addressing the conspicuous absence of holistic evaluations for prevalent DA techniques, we carried out an all-encompassing empirical assessment, in which more than 15 DA strategies were subjected to scrutiny across 8 UCR time-series datasets, employing ResNet and a multifaceted evaluation paradigm that includes accuracy, method classification and residual analysis, resulting in a benchmark accuracy of 88.94 ± 11.83%. Our investigation underscored the inconsistent efficacies of DA techniques, with certain strategies such as Permutation (boasting an average rank of 4.125 and an accuracy pinnacle of 89.20 ± 11.95%, and outperforming in 2 datasets) significantly propelling model performance, whereas others, like Rotation (84.75 ± 12.92%), inadvertently compromised it. Furthermore, we identified that the inherent attributes of datasets exert a profound impact on the success of DA, leading us to provide precise and pragmatic recommendations, derived from our empirical insights, to help practitioners align their choice of DA techniques with specific dataset nuances. To improve practical use, we’ve consolidated most of these methods into a unified Python package, whose user-friendly interface facilitates experimenting with various augmentation techniques, offering practitioners and researchers a more convenient tool for innovation than currently available options.
@@ I-A Background and Motivation
In the realm of machine learning, the quest for robust model performance often hinges on the availability of extensive datasets, a requirement that poses a significant challenge given the frequent scarcity of such data [1]. Data augmentation emerges as a pivotal solution in this context, ingeniously expanding limited datasets by generating additional, analogous data, thereby circumventing the constraints imposed by data insufficiency. This technique, renowned for its cost-effectiveness and efficacy, has gained immense traction across various machine learning disciplines, significantly enhancing model accuracy in environments characterized by limited data. In computational vision, for instance, the augmentation of images is relatively straightforward, with a plethora of simplistic yet potent methods capable of generating discernible images even under conditions of noise addition or partial image cropping [2].
@@
Contrastingly, the augmentation of time series data has not garnered comparable focus. Existing strategies often overlook the intrinsic properties defining time series data, such as temporal correlation, a critical aspect that distinguishes it from other data types. The potential for transforming time series data into frequency or time-frequency domains further complicates the application of conventional augmentation techniques, particularly those borrowed from image or speech processing, leading to potential inaccuracies in synthetic data generation. Moreover, the suitability of data augmentation methods is intricately linked to the specific task, necessitating a nuanced approach that considers the distinct requirements of tasks such as forecasting versus classification within the time series context.
I-B Objectives and Contributions
This manuscript presents an exhaustive exploration into the contemporary landscape of Data Augmentation (DA) for Time Series Classification (TSC), delving into the myriad techniques and methodologies that have sculpted the domain over the preceding decade and furnishing readers with a profound comprehension of its evolution.
1. Comprehensive Survey & Novel Taxonomy: Endeavoring to impose order and coherence upon the diverse techniques permeating this field, we unveil a pioneering taxonomy tailored explicitly for DA in TSC. Our classification schema compartmentalizes methods into quintessential categories: Transformation-Based, Pattern-Based, Generative, Decomposition-Based, and Automated Data Augmentation, thereby facilitating a methodical and streamlined approach to DA for TSC study. This structured paradigm is instrumental in the comparative analysis and judicious selection of techniques, calibrated to the nuances of specific tasks.
2. Detailed Evaluation: In a move transcending the conventional survey and taxonomy, this study makes a seminal contribution by undertaking a rigorous evaluation of more than fifteen prominent DA for TSC methodologies. These methods were subjected to a battery of tests employing a proven neural network model, establishing a robust and reliable performance benchmark. Our empirical scrutiny provides an all-encompassing perspective on the relative merits and demerits of these techniques, yielding critical insights of immense value to both the scholarly and practical realms.
@@ I-B Objectives and Contributions
3. Unified Library: Complementing these efforts, we have orchestrated a Python library, encapsulating the aforementioned time series data augmentation algorithms, designed for effortless implementation. This initiative stems from the recognition of the existing tools’ non-intuitive nature, necessitating a more user-friendly alternative.
The revelations from our investigative journey contribute a panoramic view of the current techniques, simultaneously illuminating prospective avenues for refinement and future inquiry. Our aspiration is to catalyze further innovation in DA for TSC, charting the course for the emergence of increasingly sophisticated, efficient methods adept at navigating the complexities intrinsic to time series data. The principal contributions of this manuscript are encapsulated as follows:
@@
  • We orchestrate a meticulous review of the advancements characterizing the past decade in Data Augmentation for Time-Series Classification (DA for TSC);
@@
  • We execute an in-depth evaluation of over fifteen DA for TSC techniques, utilizing a renowned neural network model, and offer an exhaustive discourse on the findings;
  • We introduce a Python compendium of preeminent time series data augmentation algorithms, conceived to present a unified, user-friendly platform for the application of diverse data augmentation strategies to time series datasets.
The structure of this article is meticulously organized to facilitate a comprehensive understanding of our investigation, as depicted in Figure 1. Section 2 embarks on a thorough literature review, delineating the state of the art in Domain Adaptation (DA) for Time Series Classification (TSC), establishing the foundation upon which our study is predicated. In Section 3, we elaborate on the methodological framework adopted for our survey, including the criteria for selection and analysis of relevant works, and engage in a detailed discussion of the findings derived therefrom. Section 4 introduces a novel taxonomy for DA techniques, methodically categorizing these methodologies into five distinct principal categories, thereby offering a structured lens through which the current landscape can be viewed. Section 5 is dedicated to the elucidation of our experimental design, meticulously detailing the setup employed in our empirical study and articulating the research questions poised at the outset. Subsequently, Section 6 addresses these research questions, presenting a critical analysis of the experimental outcomes and drawing insights pertinent to the field of DA for TSC. Section 7 ventures into an exploration of the limitations inherent to our study and the potential threats to the validity of our findings, fostering a transparent discourse on the scope and applicability of our conclusions. The article culminates in Section 8, where we synthesize our contributions to the domain and outline directions for future research.
@@ III-B Paper Collection Methodology of the Survey
The “keywords” used for searching and collecting are listed below. [DA for TSC] means Data Augmentation for Time-Series Classification, which is the focus of this research project. The search strategy encompassed several iterations, as detailed below:
@@ III-C Result and Discussion of the Survey
[DA for TSC] + Survey / Review 20 15 3
[DA for TSC] + Techniques / Methods 258 92 57
[DA for TSC] + Evaluations / Comparison 70 35 16
Query - - 76
Snowball - - 26
Overall - - 102
The terms “Hits” and “Title” refer to the number of papers returned by the search, as well as the number of papers that remain after title and body framing, respectively. The search strategy yielded 102 relevant publications, with a significant proportion originating from keyword searches and the remainder from snowballing. The annual publication trends indicate a burgeoning interest in this research area, particularly post-2019.
@@ V Experimental Study Setup
In this section, we detail the comprehensive comparative assessments conducted, utilizing various data augmentation techniques. These evaluations are designed to provide an empirical comparison between different data augmentation strategies. The primary focus of the comparison is to assess the influence of these methods on the performance of neural network models, especially in the context of time series datasets. This approach allows for a deeper understanding of how these techniques can be optimized to enhance model performance, offering a broader view of their potential impact across various applications.
@@ V-A Datasets
This research encompasses the application of several time-series classification techniques to diverse datasets from the University of California, Riverside (UCR) time-series classification repository [13]. The UCR repository is a rich collection of time-series datasets, extensively utilized within the academic community for benchmarking time series classification algorithms. For this study, eight representative categories were chosen, each representing a unique type of time-series data:
• Image (Yoga): Time-series representations of images, posing challenges in capturing spatial-temporal dynamics.
• Sensor (FordB): Sensor data, often fraught with noise, embodying the challenges of real-world data collection environments.
• Simulated (CBF): Simulated datasets, offering controlled environments where specific time-series phenomena can be isolated and studied.
• Device (ScreenType): Illustrating the challenges posed by device-specific readings, which may vary based on individual device characteristics.
• Motion (GunPointAgeSpan): Encompassing the complexities of motion data and its unique temporal dynamics.
• Spectro (Strawberry): Spectrograms and related time-frequency representations, posing challenges due to their multi-modal data nature.
• ECG (ECG5000, ECG200): Medical time-series data, critical for its implications in health and diagnostics, yet complex due to its sensitivity to minor temporal shifts.
@@
The datasets selected for this study underwent preprocessing to ensure compatibility with the data augmentation techniques. In Table II, ”Class” refers to the number of distinct categories or labels within the dataset. It reflects the granularity and diversity of events or phenomena that the dataset captures. A higher number of classes generally suggests a more complex classification task. ”Length” denotes the duration of each individual time series in the dataset, representing the number of time-points or observations for each sequence. This factor is crucial as longer sequences might capture more detailed dynamics but also introduce challenges in computational complexity and potential for overfitting.
Firstly, all datasets were rescaled to ensure that the minimum value in the training set is -1 and the maximum is 1. This normalization step helps preserve relative values and maintain consistency across different datasets. Additionally, to accommodate batch training procedures requiring input data of identical dimensions, datasets of varying lengths were zero-padded. This process involved adding zeros at the end of shorter sequences to match the length of the longest sequence in the dataset. Lastly, we addressed missing values, if present, by substituting them with zero, a standard approach that maintains the dimensional integrity of the data while ensuring that missing values do not influence the learning process.
@@ V-B Evaluated Network Models
For our evaluation, we employed the deep Convolutional Neural Network (CNN) model, ResNet, known for its utilization of residual connections. The hyperparameters for the 1D ResNet used in this evaluation were informed by research conducted by Fawaz et al. [3] and Iwana et al. [4], where they noted improvements in time series classification using data augmentation.
As shown in Figure 10, our implementation of ResNet diverges from the original model [105] by incorporating only three residual blocks with varying filter lengths and omitting max pooling. The first block employs three layers of 64 1D convolutions, the second block contains three layers of 128 1D convolutions, and the third block features three layers of 128 convolutions. Each residual block has three convolutional layers with filter sizes of 8, 5, and 3, respectively. Batch normalization [106] and ReLU activation functions follow each convolution. Additionally, the residual connections merge the input of each residual block with the input of the subsequent block through an addition operation. The final two layers of ResNet comprise a Global Average Pooling (GAP) layer and an output layer with softmax activation.
All network parameters underwent initialization using Glorot’s Uniform method [107]. We trained the network using the Adam optimizer, with an initial learning rate of 0.001 and exponential decay rates for the first and second momentum estimates at 0.9 and 0.999, respectively, following Fawaz et al.’s methodology [3].
@@
Figure 10: The architectures of the neural networks used to evaluate the data augmentation methods [4]
@@
To maintain consistency in the comparison of data augmentation methods, we ensured each model underwent the same 10,000 iterations, reducing the learning rate by 0.1 on plateaus of 500 iterations. The number of patterns in the training dataset was quadrupled compared to the original dataset. The choice of 10,000 iterations stemmed from observations indicating that the training loss for the model on each dataset typically converged before this point.
@@
Moreover, as the number of iterations remained constant, the control experiment without augmentation effectively repeated the same patterns to match the augmented experiments. Regarding hyperparameters selection, we adhered to the best practices documented in literature as closely as possible, avoiding excessive adjustments to maintain a fair comparison of the data augmentation effects.

### v5

@@ Abstract
Data Augmentation (DA) has become a critical approach in Time Series Classification (TSC), primarily for its capacity to expand training datasets, enhance model robustness, introduce diversity, and reduce overfitting. However, the current landscape of DA in TSC is plagued with fragmented literature reviews, nebulous methodological taxonomies, inadequate evaluative measures, and a dearth of accessible and user-oriented tools. This study addresses these challenges through a comprehensive examination of DA methodologies within the TSC domain. Our research began with an extensive literature review spanning a decade, revealing significant gaps in existing surveys and necessitating a detailed analysis of over 100 scholarly articles to identify more than 60 distinct DA techniques. This rigorous review led to the development of a novel taxonomy tailored to the specific needs of DA in TSC, categorizing techniques into five primary categories: Transformation-Based, Pattern-Based, Generative, Decomposition-Based, and Automated Data Augmentation. This taxonomy is intended to guide researchers in selecting appropriate methods with greater clarity. In response to the lack of comprehensive evaluations of foundational DA techniques, we conducted a thorough empirical study, testing nearly 20 DA strategies across 15 diverse datasets representing all types within the UCR time-series repository. Using ResNet and LSTM architectures, we employed a multifaceted evaluation approach, including metrics such as Accuracy, Method Ranking, and Residual Analysis, resulting in a benchmark accuracy of 84.98 ± 16.41% in ResNet and 82.41 ± 18.71% in LSTM. Our investigation underscored the inconsistent efficacies of DA techniques, for instance, methods like RGWs and Random Permutation significantly improved model performance, whereas others, like EMD, were less effective. Furthermore, we found that the intrinsic characteristics of datasets significantly influence the success of DA methods, leading to targeted recommendations based on empirical evidence to help practitioners select the most suitable DA techniques for specific datasets. To improve practical use, we have consolidated most of these methods into a unified Python Library, whose user-friendly interface facilitates experimenting with various augmentation techniques, offering practitioners and researchers a more convenient tool for innovation than currently available options.
@@ I Introduction
In machine learning, the pursuit of robust model performance often depends on the availability of extensive datasets, a challenge exacerbated by the frequent scarcity of data [1]. Data augmentation addresses this issue by generating additional, analogous data to expand limited datasets, thereby mitigating data insufficiency constraints. This technique, renowned for its cost-effectiveness and efficacy, has gained immense traction in various machine learning disciplines, significantly improving model accuracy in environments characterized by limited data. For instance, in computer vision, image augmentation is relatively straightforward, with a variety of simple yet effective techniques capable of producing discernible images even when subjected to noise or partial cropping [2].
@@
In contrast, time-series data augmentation has received less attention. Existing strategies often overlook the intrinsic properties defining time-series data, such as temporal correlation, a critical aspect that distinguishes it from other data types. The potential for transforming time-series data into frequency or time-frequency domains further complicates the application of conventional augmentation techniques, particularly those borrowed from image or speech processing, leading to potential inaccuracies in synthetic data generation. Additionally, the effectiveness of data augmentation methods is highly task-dependent, requiring a nuanced approach that addresses the unique demands of tasks like forecasting and classification within the time-series context.
This paper offers an extensive examination of the current state of Data Augmentation (DA) for Time Series Classification (TSC), exploring a wide range of techniques and methodologies that have shaped the field over the past decade, providing readers with a comprehensive understanding of its evolution.
- Comprehensive Survey and Novel Taxonomy: To bring clarity and structure to the diverse array of techniques in this field, we introduce a novel taxonomy specifically designed for DA in TSC. This classification system organizes methods into five fundamental categories: Transformation-Based, Pattern-Based, Generative, Decomposition-Based, and Automated Data Augmentation, enabling a systematic and streamlined approach to selecting and analyzing DA methods for TSC.
- Detailed Evaluation: Beyond a traditional survey and taxonomy, this study makes a significant contribution by rigorously evaluating nearly twenty DA techniques for TSC. These methods were tested using two established neural network models, providing a robust performance benchmark. Our empirical analysis offers a comprehensive perspective on the relative strengths and weaknesses of these techniques, delivering valuable insights for both academia and industry.
@@
- Unified Library: To complement our efforts, we have developed a Python Library that encapsulates most of the aforementioned time-series data augmentation algorithms, designed for ease of implementation. This initiative addresses the shortcomings of existing tools, offering a more user-friendly alternative.
The insights gained from our research provide a broad view of current techniques while highlighting potential areas for improvement and future research. We aim to foster further innovation in DA for TSC, paving the way for the development of more sophisticated and effective methods tailored to the complexities of time-series data. The main contributions of this manuscript are as follows:
@@
  • We conduct a thorough review of advancements in Data Augmentation for Time Series Classification (DA for TSC) over the past decade;
@@
  • We perform an in-depth evaluation of nearly twenty DA for TSC techniques, utilizing two renowned neural network models, and offer an exhaustive discourse on the findings;
  • We introduce a Python Library of preeminent time series data augmentation algorithms, offering a unified, user-friendly platform for the application of diverse data augmentation strategies to time series datasets.
The structure of this article is meticulously organized to facilitate a comprehensive understanding of our investigation, as depicted in Figure 1. Section II embarks on a thorough literature review, delineating the state of the art in Data Augmentation (DA) for Time Series Classification (TSC), establishing the foundation upon which our study is predicated. In Section III, we elaborate on the methodological framework adopted for our survey, including the criteria for selection and analysis of relevant works, and engage in a detailed discussion of the findings derived therefrom. Section IV introduces a novel taxonomy for DA techniques, methodically categorizing these methodologies into five distinct principal categories, thereby offering a structured lens through which the current landscape can be viewed. Section V is dedicated to the elucidation of our experimental design, meticulously detailing the setup employed in our empirical study and articulating the research questions poised at the outset. Subsequently, Section VI addresses these research questions, presenting a critical analysis of the experimental outcomes and drawing insights pertinent to the field of DA for TSC. Section VII ventures into an exploration of the limitations inherent to our study and the potential threats to the validity of our findings, fostering a transparent discourse on the scope and applicability of our conclusions. The article culminates in Section VIII, where we synthesize our contributions to the domain and outline directions for future research.
@@ III-B Paper Collection Methodology of the Survey
The keywords used for searching and collecting are listed below. [DA for TSC] means Data Augmentation for Time-Series Classification, which is the focus of this research project. The search strategy encompassed several iterations, as detailed below:
@@ III-C Result and Discussion of the Survey
[DA for TSC] + Survey / Review 25 20 8
[DA for TSC] + Techniques / Methods 273 102 72
[DA for TSC] + Evaluations / Comparison 90 55 26
Query - - 106
Snowball - - 41
Overall - - 147
The terms “Key Words”, “Hits” and “Title” refer to the number of papers returned by the search, the number of papers that remain after title framing and the number of papers that remain after body framing, respectively. The search strategy yielded 147 relevant publications until 10th June, 2024, with a significant proportion originating from keyword searches and the remainder from snowballing. Trends in annual publication indicate a growing interest in this research area, particularly after 2019, as shown in Figure 2.
@@ IV-D Decomposition-Based Methods
Figure 10: Illustrative Examples of Decomposition-Based Methods (EMD) on the FordB Dataset (Generated instances are depicted in orange)
@@ V Experimental Study Setup
In this section, we detail the comprehensive comparative evaluations performed, using various data enhancement techniques. These evaluations are designed to provide an empirical comparison between different data augmentation strategies. The primary focus of the comparison is to assess the influence of these methods on the performance of neural network models, especially in the context of time series datasets. This approach allows for a deeper understanding of how these techniques can be optimized to enhance model performance, offering a broader view of their potential impact across various applications.
@@ V-A Datasets
This research encompasses the application of several time series classification techniques to various datasets from the University of California, Riverside (UCR) time series classification repository [13]. The UCR repository is a rich collection of time series datasets, extensively utilized within the academic community for benchmarking time series classification algorithms. For this study, fifteen representative categories were chosen, each representing a unique type of time series data. By including these datasets, we cover all fifteen types of classification datasets within the UCR repository, namely Device, ECG, EOG, EPG, Hemodynamics, HRM, Image, Motion, Power, Sensor, Simulated, Spectro, Spectrum, Traffic, and Trajectory. These types encompass a diverse range of domains, such as medical signals, motion capture, sensor readings, and simulated environments, among others.
The UCR repository is widely recognized for its extensive collection of time-series datasets, each belonging to a distinct category that reflects different real-world applications. This ensures that our selections are not only classic and representative but also demonstrate the broad applicability and generalizability of our experiments across diverse domains. Our goal was to delve deeper into the nuanced relationships between DA methods and datasets. Selecting fewer, but highly representative datasets, allows us to achieve this objective more effectively.
@@
EOGHorizontalSignal EOG 362 362 12 1250
Fungi HRM 18 186 18 201
GestureMidAirD1 Trajectory 208 130 26 Vary
InsectEPGRegularTrain EPG 62 249 3 601
MelbournePedestrian Traffic 1194 2439 10 24
PigCVP Hemodynamics 104 208 52 2000
PowerCons Power 180 180 2 144
SemgHandMovementCh2 Spectrum 450 450 6 1500
In Table II, the term “Class” refers to the number of distinct categories or labels within the dataset. It reflects the granularity and diversity of events or phenomena that the dataset captures. A higher number of classes generally suggests a more complex classification task. “Length” denotes the duration of each individual time series in the dataset, representing the number of time points or observations for each sequence. This factor is crucial as longer sequences might capture more detailed dynamics but also introduce challenges in computational complexity and potential for overfitting.
All datasets were firstly rescaled to ensure that the minimum value in the training set is -1 and the maximum is 1. This normalization step helps preserve relative values and maintain consistency across different datasets. Additionally, we address missing values, if present, by substituting them with zero, a standard approach that maintains the dimensional integrity of the data while ensuring that missing values do not influence the learning process.
@@ V-B Evaluated Network Models
For our evaluation, we use two deep learning models: the deep convolutional neural network (CNN) model ResNet and the Long Short-Term Memory (LSTM) model. ResNet is particularly well-suited for time series data due to its ability to capture hierarchical features through convolutional layers and its robustness to overfitting via residual connections. LSTM, on the other hand, excels at learning temporal dependencies and capturing long-term trends in sequential data due to its unique memory cell architecture, which can retain information over long periods. Both ResNet and LSTM have been widely adopted in the field of data augmentation for time series classification, serving as classic models with extensive usage and validation.
Many studies have demonstrated their superior performance compared to other models, highlighting their effectiveness in this task. In our survey, over one-third of the papers utilized ResNet, LSTM, or both to evaluate their proposed algorithms, including studies such as [39, 55, 30, 3, 49, 36, 105, 50, 106, 107]. This widespread use underscores the robustness and reliability of these models in benchmarking DA methods for TSC. Our decision to use both ResNet and LSTM stems from their prevalent usage and proven success in many renowned studies. Additionally, the unique characteristics of these models make them particularly suited for time series classification. By employing these models, we aim to provide a reliable evaluation of DA methods and are confident they will demonstrate excellent results with the methods we have collected.
@@
Hyperparameters for the 1D ResNet used in this evaluation were informed by research conducted by Fawaz et al. [3] and Iwana et al. [4], where they noted improvements in time series classification using data augmentation. Similarly, the LSTM configurations were guided by studies [4, 108, 109] that have shown its effectiveness in handling temporal dependencies in time series data.
V-B1 LSTM
As shown in Figure 11, our implementation of the LSTM network is based on the architecture proposed by Karim et al. [109]. It incorporates both LSTM and convolutional layers to capture temporal dependencies and extract meaningful features from the time-series data. Although this setup uses LSTM in a non-standard way, it has been shown to be more effective than traditional configurations [4, 110].
The LSTM component begins by permuting the input shape to align with the multivariate time-step structure required by the LSTM layer. This layer consists of 128 units, followed by a dropout layer with a dropout rate of 0.8, which helps prevent overfitting and enhances generalization. The convolutional component consists of three 1D convolutional layers. The first layer applies 128 filters with a kernel size of 8 and uses the He uniform initializer for weights. The second layer increases the number of filters to 256, with a kernel size of 5. The third layer reduces the filter count to 128, using a kernel size of 3. Each convolutional layer is followed by batch normalization [111] and ReLU activation functions, ensuring stable training and promoting non-linearity. A Global Average Pooling (GAP) layer is used to flatten the output of the convolutional layers, and the results are then concatenated with the output from the LSTM layer. This combination allows the model to leverage both sequential information from the LSTM and spatial features extracted by the convolutional layers. The final output layer uses softmax activation to produce class probabilities, making the network suitable for multi-class time-series classification tasks.
@@ V-B1 LSTM
Figure 11: The architectures of the neural networks (LSTM Fully Convolutional Network) used to evaluate the data augmentation methods
V-B2 ResNet
As shown in Figure 12, our implementation of ResNet diverges from the original model [112] by incorporating only three residual blocks with varying filter lengths and omitting max pooling. The first block employs three layers of 64 1D convolutions, the second block contains three layers of 128 1D convolutions, and the third block features three layers of 128 convolutions. Each residual block has three convolutional layers with filter sizes of 8, 5, and 3, respectively. The batch normalization functions [111] and the ReLU activation functions follow each convolution. Additionally, the residual connections merge the input of each residual block with the input of the subsequent block through an addition operation. The final two layers of ResNet comprise a Global Average Pooling (GAP) layer and an output layer with softmax activation.
Figure 12: The architectures of the neural networks (ResNet) used to evaluate the data augmentation methods
@@ V-B2 ResNet
All network parameters were initialized using the Glorot uniform method [113]. We trained the network using the Adam optimizer [114], with an initial learning rate of 0.001 and exponential decay rates for the first and second momentum estimates at 0.9 and 0.999, respectively, following Fawaz et al.’s methodology [3]. To maintain consistency in the comparison of data augmentation methods, we ensured that each model underwent the same 10,000 iterations, reducing the learning rate by 0.1 on plateaus of 500 iterations. The number of patterns in the training dataset was quadrupled compared to the original dataset. The choice of 10,000 iterations stemmed from observations that indicated that the training loss for the model on each dataset typically converged before this point.
@@
Moreover, as the number of iterations remained constant, the control experiment without augmentation effectively repeated the same patterns to match the augmented experiments. Regarding hyperparameter selection, we adhered to the best practices documented in the literature as closely as possible, avoiding excessive adjustments to maintain a fair comparison of the data augmentation effects.